A measure of how hot weather feels to an average person is the Apparent Temperature Index.

A model for this index is:


A = 0.885t − 22.4h + 1.20th − 0.544,


where A is the apparent temperature in degrees Celsius, t is the air temperature, and h is the

relative humidity in decimal form.

a) Find ∂A/∂t and ∂A/∂h when t = 300 and h = 0.80.


b) Which has a greater effect on A, air temperature or humidity? Explain.


1
Expert's answer
2022-03-02T13:10:28-0500

a)


"\partial A/\partial t=0.885+1.20h"

"\partial A/\partial h=-22.4+1.20t"

"t=300, h=0.80"


"\partial A/\partial t=0.885+1.20(0.80)=1.845"

"\partial A/\partial h=-22.4+1.20(300)=337.6"

b)

Since "|\partial A/\partial h|>|\partial A/\partial t|," then relative humidity has a a greater effect on A, than air temperature.